-6|5-2y|= -9
From a homework worksheet.

Answer by scott8148(6628)

dividing by -6 gives |5-2y|=3/2 ... 5-2y=3/2 ... -2y=-7/2 ... y=7/4

and 5-2y=-3/2 ... -2y=-13/2 ... y=13/4